OpenStudy (anonymous):

A floor 5m by 20m is covered by square tiles of side 20cm. how many tiles are there?

OpenStudy (anonymous):

5*20=40 tile :P

OpenStudy (jamesj):

What's the area of the floor in square meters? Call that A

OpenStudy (jamesj):

Then you need to know the area of a tile in square meters, call that T. Then the number of tiles you need is A/T.

OpenStudy (anonymous):

100m^2/4m^2????

OpenStudy (jamesj):

I agree A = 100 m^2. Recalculate T.

OpenStudy (jamesj):

The side of a tile is 0.2 m. Hence T = ...

OpenStudy (anonymous):

00.4

OpenStudy (jamesj):

T = (0.2 m)*(0.2 m) = (0.2)^2 m^2 = ...

OpenStudy (jamesj):

Alternatively, the length of the side of a tile is 1/5 m, as 1/5 = 0.2. Hence the area of a tile is \[ T = \left( \frac{1}{5} m \right)^2 = \frac{1}{5^2} m^2 = ... what? \]

OpenStudy (anonymous):

lol i really suck

OpenStudy (jamesj):

@hc, your answer above is almost right; the decimal point is in the wrong place.

OpenStudy (anonymous):

100m^2/00.4m^2

OpenStudy (anonymous):

2500

OpenStudy (jamesj):

Right, but note: the area of a tile is 0.04 m^2, not 00.4 ... ... because 00.4 = 0.4

OpenStudy (anonymous):

thanks james